OpenStudy (anonymous):

A survey was conducted among 400 students of age groups 7–12 years and 13–18 years to find their favorite music genre. The students had to select any one genre out of jazz, rock, and pop. Out of the 200 students in the age group 7–12 years who participated in the survey, 142 liked rock or pop music. The total number of students of both age groups who liked jazz was 72. Using a two-way table, compute the total number of students in the age group 13–18 years who liked rock or pop music. 122 130 142 186

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

how

Vocaloid (vocaloid):

whoops, small correction, the circled value should be this one

Vocaloid (vocaloid):

anyway, the first row tells us that (number of kids aged 7-12 who like jazz) + (number of kids aged 7-12 who like rock or pop) = 200 find the number of kids aged 7-12 who like jazz

Vocaloid (vocaloid):

 I have to get going, but maybe this will help explain it a bit better?

Vocaloid (vocaloid):

A + 142 = 200 A + B = 72 B + C = D 200 + D = 400 the value we want to find is C

Vocaloid (vocaloid):

first find A, then B, then D, and then you can use that information to find C

OpenStudy (lynfran):

good explanation @Vocaloid

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

@Vocaloid

OpenStudy (lynfran):

to find A A+142=200 A=200-142 A=?

OpenStudy (anonymous):

58

OpenStudy (lynfran):

good now A+B=72 A=58 so 58+B=72 B=72-58

OpenStudy (anonymous):

14

OpenStudy (lynfran):

good B+C= D lets look at the last equation 200+D=400 D=400-200 D=?

OpenStudy (anonymous):

200

OpenStudy (lynfran):

good now B+C=D and B=14 and D=200 so 14+C=200 c=200-14 c=??

OpenStudy (anonymous):

186

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

thx

OpenStudy (lynfran):

welcome thanks to @Vocaloid